Scaleway
Monitoring - We're enhancing the internal routing policies in DC5.

This will allow us to offer you a better (fewer) latency between:
* fr-par-2 (dc5) baremetal services (dedibox,Elastic Metal).
* fr-par-2 (dc5) virtualized services (Instances and higher level services).

The changes will be operated on DC5 aggregators from 4th to 6th October,
no traffic loss is expected.

Thank you for your patience and cooperation!

Oct 04, 2022 - 15:07 CEST
Monitoring - A fix has been implemented and our team is monitoring the situation.
Sep 28, 2022 - 16:42 CEST
Investigating - We are experiencing an anomaly on the uplink of some Elastic Metal racks in PAR2, which causes intermittent packet losses.

Our team is already aware of the anomaly and is working to resolve it as soon as possible.

Thank you in advance for your cooperation and patience with this situation

Sep 28, 2022 - 16:15 CEST
Investigating - We are currently investigating on this issue, gmail detects our mails in spam

We work to solve this as soon as we can and this status will be updated at every iteration

Sep 21, 2022 - 13:44 CEST
Monitoring - 100% of the impacted clusters (181 in total) are now back and up and running.
The maintenance team are now monitoring all the activities.

Sep 09, 2022 - 19:14 CEST
Update - 30% of the clusters have been fully restored to nominal state (31), teh remaining, on a different etcd cluster are currently being synced (66% so far), and should be done by 7 PM Paris time.
Sep 09, 2022 - 18:08 CEST
Identified - 30% of the impacted clusters are restored, the other part is still in progress
Sep 09, 2022 - 16:44 CEST
Update - We are continuing to investigate this issue.
Sep 09, 2022 - 11:01 CEST
Investigating - Some error will be "database space exceeded" related to your Kapsule cluster.
Our team is aware of the issue and working on a fix.

Sep 08, 2022 - 21:00 CEST
Elements - AZ Under Maintenance
90 days ago
99.66 % uptime
Today
fr-par-1 Operational
90 days ago
100.0 % uptime
Today
fr-par-2 Under Maintenance
90 days ago
98.01 % uptime
Today
fr-par-3 Operational
90 days ago
100.0 % uptime
Today
nl-ams-1 Operational
90 days ago
100.0 % uptime
Today
pl-waw-1 Operational
90 days ago
99.98 % uptime
Today
nl-ams-2 Operational
90 days ago
100.0 % uptime
Today
Elements - Products Degraded Performance
90 days ago
99.5 % uptime
Today
Instances Operational
90 days ago
99.93 % uptime
Today
BMaaS Operational
90 days ago
100.0 % uptime
Today
Object Storage Operational
90 days ago
100.0 % uptime
Today
C14 Cold Storage Operational
90 days ago
99.86 % uptime
Today
Kapsule Operational
90 days ago
98.79 % uptime
Today
DBaaS Operational
90 days ago
99.98 % uptime
Today
LBaaS Operational
90 days ago
100.0 % uptime
Today
Container Registry Operational
90 days ago
99.99 % uptime
Today
Domains Operational
90 days ago
100.0 % uptime
Today
Elements Console Operational
90 days ago
94.87 % uptime
Today
IoT Hub Operational
90 days ago
99.99 % uptime
Today
Account API Operational
90 days ago
99.99 % uptime
Today
Billing API Degraded Performance
90 days ago
100.0 % uptime
Today
Functions and Containers Operational
90 days ago
99.76 % uptime
Today
Block Storage Operational
90 days ago
100.0 % uptime
Today
Elastic Metal Operational
90 days ago
98.01 % uptime
Today
Apple Silicon M1 Operational
90 days ago
100.0 % uptime
Today
Private Network Operational
90 days ago
100.0 % uptime
Today
Dedibox - Datacenters Partial Outage
90 days ago
99.15 % uptime
Today
DC2 Operational
90 days ago
98.94 % uptime
Today
DC3 Operational
90 days ago
99.94 % uptime
Today
DC5 Partial Outage
90 days ago
97.72 % uptime
Today
AMS Operational
90 days ago
100.0 % uptime
Today
Dedibox - Products Operational
90 days ago
98.76 % uptime
Today
Dedibox Operational
90 days ago
99.38 % uptime
Today
Hosting Operational
90 days ago
92.03 % uptime
Today
SAN Operational
90 days ago
98.74 % uptime
Today
Dedirack Operational
90 days ago
100.0 % uptime
Today
Dedibackup Operational
90 days ago
100.0 % uptime
Today
Dedibox Console Operational
90 days ago
99.99 % uptime
Today
Domains Operational
90 days ago
100.0 % uptime
Today
RPN Operational
90 days ago
99.96 % uptime
Today
Miscellaneous Operational
90 days ago
100.0 % uptime
Today
Excellence Operational
90 days ago
100.0 % uptime
Today
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Major outage
Partial outage
No downtime recorded on this day.
No data exists for this day.
had a major outage.
had a partial outage.
Scheduled Maintenance
Temporary ticket history unavailability Oct 10, 2022 10:00-13:00 CEST
Following an internal maintenance on October 10th, tickets created before October 10th 10am CEST will not be available on customers' console during the maintenance window.
Maintenance should last for 3 hours.

New tickets' creation will still be available during this timeframe.

Posted on Oct 03, 2022 - 17:12 CEST
[DC5] Maintenance of inverters Oct 20, 2022 08:00-17:00 CEST
As part of the curative maintenance, we will going to do maintenance of the inverters of the electrical chains of the TGBT 7.1 and 7.2
The inverters of these electrical chains supply the bays in computer rooms 1 and 2.

The operations will be conducted by our maintainer and under our permanent supervision.
This maintenance requires isolating each inverter in order to work on the device.

The operation will therefore take place inverter by inverter and will have no impact on your equipment hosted in our Datacenter: the other inverters in the electrical chain are sized so as to recover the load from the inverter when stopped.

Posted on Sep 28, 2022 - 16:03 CEST
As part of the preventive maintenance operations of the Datacenter, we will carry out mechanical maintenance of the
generators followed by an infrastructure redundancy test.

The operation consists of simulating a power cut to the TGBTs and validating load recovery as well as operation on
generator sets after électrical maintenance of each set according to the following provisional schedule:

• GE LT1.1: 21/11/22 (between 8:30 a.m. and 4:00 p.m.)
• GE LT1.2: 22/11/22 (between 8:30 a.m. and 4:00 p.m.)
• GE LT2.1: 23/11/22 (between 8:30 a.m. and 4:00 p.m.)
• GE LT2.2: 24/11/22 (between 8:30 a.m. and 4:00 p.m.)
• GE BAT: 25/11/22 (between 8:30 a.m. and 4:00 p.m.) No impact for IT.
• GE LT7.1: 28/11/22 (between 8:30 a.m. and 4:00 p.m.)
• GE LT7.2: 29/11/22 (between 8 a.m. and 2 p.m.)

Our teams as well as the usual maintenance companies will be mobilized for the smooth running of operations.
This intervention does not present any particular risk of service interruption and will be processed on worked hours.

Posted on Sep 23, 2022 - 12:26 CEST
Past Incidents
Oct 4, 2022
Resolved - Our team finished to apply the update and fix issue
Oct 4, 18:04 CEST
Identified - Our team is still working on this issue
Sep 28, 12:27 CEST
Investigating - We're facing issues on email transfers, this means that some mail transfers are not working

We work to solve this as soon as we can and this status will be updated at every iteration

Sep 21, 13:40 CEST
Resolved - Issue has been resolved, API keys creation is working properly again.
Oct 4, 11:36 CEST
Investigating - Our teams have noticed an internal outage currently making it not possible to create new API keys.

Issue is being investigated, we will update this status as soon as we have additional details.

Oct 4, 11:32 CEST
Oct 3, 2022
Resolved - The incident has now been resolved and the situation was being monitored all morning.
Our helpdesk is now responding as expected.

We thank you for your patience and cooperation!

Oct 3, 17:28 CEST
Investigating - We are currently updating our system, you might be facing some delays on replies for all support related requests.
We will update this status as soon as the situation evolves

Oct 2, 10:58 CEST
Resolved - Due to a crash of one of the RPN routers on our level during a maintenance check, the incident lasted 7 minutes:

Incident start : 09:57 UTC
Incident end : 10:04 UTC

The RPN connectivity of the following racks was cut in Room 1 of our DC5:

C36, C39, C41, C42, C50, C51, C54, C55, C56, C61, C62, C64, C65, C69, D39, D40, D48, D49, D50, D51, D52, D54, D55, D58, D59, D60, D61, D62, D63.

The public connection was not impacted by this incident.

Oct 3, 10:00 CEST
Oct 2, 2022
Oct 1, 2022

No incidents reported.

Sep 30, 2022

No incidents reported.

Sep 29, 2022
Resolved - This incident has been resolved.
Sep 29, 11:51 CEST
Update - The incident started this morning at 8AM
Sep 29, 11:32 CEST
Investigating - We are currently facing an error on one of our etcd server.
Our team is aware of the issue and working on a fix.

Sep 29, 08:00 CEST
Sep 28, 2022
Resolved - The incident has been resolved by our team in charge.

If the situation continues at your level, we invite you to create a trouble ticket with our support team if needed.

We thank you for your patience and cooperation in this situation.

Sep 28, 14:05 CEST
Investigating - We have detected an issue on san-ssd-2.rpn.online.net
This RPN-SAN is currently unavailable.

Issue has been forwarded to our team for resolution.

Sep 28, 11:49 CEST
Sep 27, 2022
Resolved - This incident has been resolved.
Sep 27, 12:08 CEST
Update - We are currently re-importing the database to fix the issue.
We will update the status when the fix will be deployed.

Sep 27, 11:00 CEST
Identified - The anomaly has been identified by our team, a fix will be deployed quickly to resolve the situation.

We thank you for your patience and cooperation!

Sep 26, 09:39 CEST
Investigating - We are currently investigating on this issue

We work to solve this as soon as we can and this status will be updated at every iteration

Sep 25, 11:07 CEST
Sep 26, 2022
Resolved - The problem is now solved. Feel free to reach our support team if need.
Thank you for your patience and cooperation!

Sep 26, 11:30 CEST
Monitoring - A fix has been implemented and we are monitoring the results.
Sep 23, 20:41 CEST
Identified - Our teams have identified an issue that may cause slowness on actions on your DBaaS in the PAR1 location, including creation, autohealing or upgrades.
Sep 23, 19:00 CEST
Resolved - The problem is now solved. Feel free to reach our support team if need.
Sep 26, 11:15 CEST
Monitoring - The service is back online and we are currently monitoring the server.
Sep 23, 16:46 CEST
Investigating - We are currently facing an error on one of our etcd server.
Our team is aware of the issue and working on a fix.

Sep 23, 15:00 CEST
Sep 25, 2022
Resolved - The public switch has been replaced and reconfigured.
It is now up and running and public connectivity is available again on the rack.

Sep 25, 12:52 CEST
Investigating - We are currently encountering an issue in DC3 Room 4 4-5 Rack F15.
The public switch of this rack is currently down, servers have no public connectivity.

Sep 25, 10:14 CEST
Resolved - The maintenance was completed on September 19th at 7pm.
Thank you for your cooperation!

Sep 25, 10:33 CEST
Identified - A maintenance is in progress on a transformer at DC3. This will last all day.
Sep 19, 12:42 CEST
Resolved - This incident has been resolved.
Sep 25, 10:12 CEST
Monitoring - The issue seems fix for now, do not hesitate to come back to the ticket if you are still experiencing the error.
Sep 21, 14:07 CEST
Investigating - We have detected a switch down in DC2 Room 103 Rack J3
Servers in that rack currently have no public network access and are unreachable.
Issue has been forwarded to our team for resolution.

Sep 21, 05:56 CEST
Sep 24, 2022

No incidents reported.

Sep 23, 2022
Resolved - The issue has been resolved
Sep 23, 14:18 CEST
Investigating - The SSL certificate cannot be generated.
Our team is looking into it

Sep 20, 14:20 CEST
Resolved - The incident has been resolved
Sep 23, 13:17 CEST
Investigating - We are facing an issue using php 7.3 on platform 14, this means that this version of php on this platform return a file to download.

Alternatively, you could use php 7.2

We work to solve this as soon as we can and this status will be updated at every iteration

Sep 21, 14:17 CEST
Resolved - This incident has been resolved.
Sep 23, 12:36 CEST
Investigating - Our team is investigating on an issue related to an error on email setting from te webmail.
Sep 21, 11:07 CEST
Resolved - The incident has been resolved
Sep 23, 11:15 CEST
Investigating - We are currently investigating this issue.
Our team is actively working to resolve this incident.

Sep 23, 08:59 CEST
Completed - The scheduled maintenance has been completed.
Sep 23, 09:00 CEST
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Sep 19, 09:01 CEST
Scheduled - Following the commissioning of a new bar duct, a migration of one of the two electrical channels is required on the servers in production in room 7/C.

Impact: Each rack listed below will be powered by a single channel for approximately 1⁄2 hour per rack.
Room 7 - Bay C1 to C16 - 19/09/2022
Room 7 - Bay C17 to C30 - 20/09/2022
Room 7 - Bay C31 to C45 - 21/09/2022
Room 7 - Bay C46 to C60 - 22/09/2022
Room 7 - Bay C61 to C73 - 23/09/2022

Aug 25, 15:00 CEST
Sep 22, 2022
Resolved - This incident has been resolved.
Sep 22, 17:57 CEST
Investigating - Customer on cloud hosting platform cannot access their dashboard
Sep 22, 15:59 CEST
Resolved - This incident has been resolved.
Sep 22, 14:50 CEST
Identified - The product database as a service has encountered some issue, some of them were unavailable this morning Since 11:09 to 11:25 am.
Sep 22, 11:38 CEST
Sep 21, 2022
Resolved - This incident has been resolved.
Sep 21, 16:45 CEST
Monitoring - The issue seems fixed now, our team is still monitoring the service.
Sep 21, 14:28 CEST
Investigating - Our services are investigating an incident related to database updates
They can randomly get stuck during the update

Sep 21, 08:26 CEST
Resolved - This incident has been resolved.
Sep 21, 08:04 CEST
Investigating - We're facing issues to unlist our SMTP servers from few RBL listing, that means some of your mail may not reach your recipient as expected.

We work to solve this as soon as we can and this status will be updated at every iteration

Sep 12, 09:39 CEST
Resolved - This incident has been resolved.
Sep 21, 08:03 CEST
Investigating - We are currently encountering a problem sending random emails on the webhosting classic service

We work to solve this as soon as we can and this status will be updated at every iterations

Sep 10, 09:26 CEST
Resolved - The incident has been fixed
Sep 21, 08:02 CEST
Investigating - Our teams have identified an issue regarding automatic SSL certificate generation.
As a result, HTTPS protocol is not usable if your certificate requires to get renewed.

We will update this status as soon as we have additional details.

Sep 9, 16:36 CEST
Sep 20, 2022
Resolved - The incident has been solved and it back to normal.
Sep 20, 10:13 CEST
Investigating - Our site https://webmail.online.net is not currently accessible. The team in charge has been notified and is investigating this issue. We will update this status as soon as we have news.
Sep 19, 22:00 CEST